3. Bricks are delivered to a work site and stacked in rows and columns, forming a rectangular prism. The length of the prism is 1 foot greater than its width, and its height is 2 feet less than its width. Find the dimensions of the prism formed by the bricks, given that its volume is 40 cubic feet.
Answer by ikleyn(52797)

.
Length * Width * Height = 40


L      * W     * H      = 40


L*H*W = 40


(W+1)*W*(W-2) = 40


  5  *4*2     = 40


W = 4,  L = 5,  H = 2.